OpenStudy (anonymous):

(PLEASE HELP) River C is 100 miles longer than RIver D. If the sum of their length is 5,500 miles what is the length of each river?

OpenStudy (chaise):

Let C equal the length of river C Let D equal the length of river D C+D=5500 C+100=D Do you need more help solving this set of simultaneous equations?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

here is one

OpenStudy (chaise):

C+C+100=5500 2C+100=5500 2C=5400 C=2700 2700+D=5500 D=5500-2700 D=2800

OpenStudy (anonymous):

ah thanks!

OpenStudy (anonymous):

a. Solve the system of linear equations using substitution. x = 8y - 4 [03:37:28] Nicoluu-Chan<3: x + 8y = 6 [03:38:13] Nicoluu-Chan<3: b. Which expression would be easier to substitute into the other equation, in order to solve the problem? explain your reasoning

OpenStudy (anonymous):

oh